OpenStudy (anonymous):

(2+3i)x + (3+2i)y =1

OpenStudy (itsonlycdeee):

What are you supposed to be looking for? The equation of a circle?

OpenStudy (anonymous):

I need to find x-y

OpenStudy (anonymous):

x-y=?

OpenStudy (itsonlycdeee):

so then you need to find a common denominator right?

OpenStudy (anonymous):

yeah, i dont understand imaginary numbers very much.

OpenStudy (anonymous):

Tryed to square it, but, that doesnt seem to work

OpenStudy (anonymous):

Lets see, after multiplying i get 2x + 3xi + 3y + 2iy =1

OpenStudy (anonymous):

I can factorise it, but again, i dont understand imaginary numbers much, and cant work out this problem

jimthompson5910 (jim_thompson5910):

(2+3i)x + (3+2i)y =1 2x+3i*x + 3y+2i*y =1 (2x+3y)+(3i*x+2i*y) =1 (2x+3y)+(3i*x+2i*y) =1 + 0i (2x+3y)+(3x+2y)i =1 + 0i so this means 2x+3y = 1 3x+2y = 0

OpenStudy (anonymous):

Tyvm :D